11 0 10 0 1 1 0 Find the eigenvalues and the corresponding eigenvectors of the matrix 0 0 2 0 Lo 0 0 2][عدل]
Answer #1

Given Matrix 0 - O 2 0 o Eigen valuver characteristic Equation = |A-AB120 O O O O O 0 (1-d) Cird) (2-d) (2-d) ao 1 d=lor da llet be the eigen vector 3* Jr Solving (A-di] (n) = we get in. u y z >>> y=0 2-0 18) 41] **8 w=x=1 then 11 and 16 are the eiĪT way X=4 J ::22 [ ] and are the eigen vectors for 1:2 Therefore, given matrix has eigen valur, A1,1,2,2 and the eigen avect